# Closing the intent-to-behavior gap via Fulfillment Priority Logic

### v2

@@ III-D Fulfillment Q-values (FQ-values)
$\mathbf{FQ}^{\pi}(s,a)=(1-\gamma)\mathbf{Q}^{\pi}(s,a)\in[0,1]^{n}\quad\forall s\in S,a\in A\vspace{-0.7em}$ (2)
@@ IV-A Power Mean Foundations
[…]
@@
Monotonicity in $\text{Monotonicity in }p\quad$ $p_{1}<p_{2}\implies\overline{\mu}_{p_{1}}(\vec{\mathbf{x}})\leq\overline{\mu}_{p_{2}}(\vec{\mathbf{x}})$
@@ IV-B Logical Language
  • $[\phi]_{\delta}$ offsets the priority of $\phi$ by $\delta\in[-1,1]$.
@@
[…]
$u(\phi_{1}\land^{p}\phi_{2})$ $:=\overline{\mu}_{p}(u(\phi_{1}),u(\phi_{2}))$ for $\text{for }\phi_{1},\phi_{2}:\text{{FPL}}$ (4)
$u(\neg\phi)$ $:=1-u(\phi)$ for $\text{for }\phi:\text{{FPL}}$ (5)
$u(\phi_{1}\lor^{p}\phi_{2})$ $:=u(\neg(\neg\phi_{1}\land^{p}\neg\phi_{2}))$ for $\text{for }\phi_{1},\phi_{2}:\text{{FPL}}$ (6)
$u([\phi]_{\delta})$ $:=\frac{u(\phi)+\max(\delta,0)}{1+\delta}$ for $\text{for }\phi:\text{{FPL}}$ (7)
These semantics preserve logical relationships between objectives. The conjunction $\land^{p}$ combines objectives that must be fulfilled together, with $p\leq 0$ controlling strictness (formalized in Definition IV-G). The disjunction $\lor^{p}$, defined via De Morgan’s laws, allows focusing on the objective with highest marginal benefit. The offset $[\phi]_{\delta}$ creates priority by raising the baseline fulfillment of $\phi$, implementing a form of lexicographic ordering where objectives with lower offsets are prioritized. By making fulfillment composition explicit, FPL allows practitioners to seperate concerns and reason about the trade-offs between objectives in a principled way. We expect practitioners to first define fulfillment values for their objectives, testing whether their intention for the fulfillment of objectives maps well on to [0,1], and then use FPL to compose them in a way that is consistent with their priorities.
@@ IV-C Relation to Fuzzy Logic
FPL similar to fuzzy logic, generalizes boolean operations to the continuous domain $[0,1]$, meaning, at the limits $\{0,1\}$ it is equivalent to boolean logic. Choosing $p=0$, our conjunction operator becomes the geometric mean ($\forall_{x,y\in[0,1]},x\land^{0}y=\sqrt{x\cdot y}$) closely resembling the product t-norm [41] ($\forall_{x,y\in[0,1]},x\land y=x\cdot y$) commonly used in fuzzy logic. Also, when $p\to-\infty$, $\land^{p}$ becomes the minimum operator, which is equivalent to the fuzzy logic’s minimum t-norm.
@@ IV-E Toy Example
Priority Offset ($[{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}f_{0}}]_{\delta}\land^{p}{\color[rgb]{0.1171875,0.56640625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.56640625,1}f_{1}}$) [Uncaptioned image]: The offset gives precedence to ${\color[rgb]{0.1171875,0.56640625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.56640625,1}f_{1}}$ until it reaches the threshold established by $\delta$, at which point ${\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}f_{0}}$ becomes important as well. This priority-based curriculum allows for coherent optimization, as opposed to methods that change the reward function requiring Q-values to be re-trained.
@@ IV-G Guarantees in FPL
  This bound guarantees that when a power mean outputs value $y$, every input component must have at least fulfillment ${}^{p}]{n(y^{p}-1)+1}$.
@@
This bound is crucial for practitioners providing concrete guarantees about objective fulfillment. For example, with two objectives ($n=2$) and $p=-2$, achieving an output of 0.9 guarantees each individual objective has at least 0.38 fulfillment. This minimum-fulfillment is particularly important in safety-critical applications where we need to ensure no objective is severely underperforming. The parameter $p$ allows practitioners to trade off between stronger guarantees (more negative $p$) and easier optimization (less negative $p$), as stricter bounds require more precise balancing of objectives.
@@
  where $\mathbbm{1}_{i}$ denotes a vector of zeros with a 1 in position $i$. This states that for any change $\delta$ to component $i$, there exists a change $\delta^{\prime}$ to component $j$ that maintains the same power mean.
@@
  For any vector $\vec{\mathbf{f}}$, there exists a vector with $n-1$ ones and a value $v$ that has the same power mean but with $v$ being bounded from above by the minimum of $\vec{\mathbf{f}}$.
@@
For a vector with $n-1$ ones and one value $f$, if $\overline{\mu}_{p}((\vec{\mathbf{1}}_{n-1},f))=y$ then:
@@ V-A Mitigating Overestimation Bias
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{FV}^{\text{obs}}=(1-\gamma)\sum_{t=0}^{n-1}\gamma^{t}\vec{\mathbf{r}}_{t}+\textsc{truncated}\cdot(1-\gamma)\vec{\mathbf{r}}_{n}\frac{\gamma^{n}}{1-\gamma}}\vspace{-0.5em}$
@@
[…]
@@ VI-C1 Pendulum-v1
$-\theta^{2}-0.1\,\dot{\theta}^{2}-0.001\,\text{torque}^{2}$ $F_{\text{angle}}^{2}\land^{p}F_{\text{actuation}}$
@@ VI-C2 Reacher-v4
$-\text{distance}-0.1||torque||^{2}$ $F_{\text{distance}}^{2}\land^{p}(\vec{\mathbf{F}}_{\text{torque}})^{\land p}$
@@ VI-C3 Hopper-v4
$1+\frac{dx}{dt}-0.001\cdot||action||^{2}_{2}$ $(\vec{\mathbf{F}}_{\text{speed}})^{\land p}\land^{p}(\vec{\mathbf{F}}_{\text{action}})^{\land p}$
@@ VI-C4 LunarLanderContinuous-v2
[…]
@@
Here the offsets create a natural curriculum during training: the agent first focuses on basic proximity ($F_{\text{near}}$), then simultaneously addresses precise positioning, leg contact, and landing ($[F_{\text{very\_near}}]_{0.1}$, $[F_{\text{legs}}]_{0.1}$, $[F_{\text{landed}}]_{0.1}$), and finally optimizes fuel efficiency ($[F_{\text{fuel}}]_{0.5}$) once the primary landing objectives are reasonably satisfied. The conjunction ensures all objectives must ultimately be satisfied for successful landing.
@@ VI-D Behavioral Analysis
A note on parameter selection in FPL: FPL is robust to reasonable variations in power mean parameters and offsets, we choose $p$ as either 0 or $-1$, which primarily serve to optimize sample efficiency rather than fundamentally changing the desired behavior. For example, not squaring the angle term in Pendulum would still result in an upright pendulum, but with slower convergence due to more conservative actions. This behavioral consistency persists across training runs, unlike linear weighted reward functions that often converge to different local optima depending on initialization.

### v3

@@ III-D Fulfillment Q-values (FQ-values)
$\mathbf{FQ}^{\pi}(s,a)=(1-\gamma)\mathbf{Q}^{\pi}(s,a)\in[0,1]^{n}\quad\forall s\in S,a\in A\vskip-6.99997pt$ (2)
@@ IV-A Power Mean Foundations
$p\to$-$\infty$ $:\text{Minimum}$ $p=1$ $:\text{Arithmetic mean}$
$p=-1$ $:\text{Harmonic mean}$ $p=2$ $:\text{Root mean square}$
$p=0$ $:\text{Geometric mean}$ $p\to\infty$ $:\text{Maximum}$
@@
$\text{Monotonicity in }p\quad$ $p_{1}<p_{2}\implies\overline{\mu}_{p_{1}}(\vec{\mathbf{x}})\leq\overline{\mu}_{p_{2}}(\vec{\mathbf{x}})$
@@ IV-B Logical Language
  • $\phi]_{\delta}$ offsets the priority of $\phi$ by $\delta\in[-1,1]$.
@@
$u(f)$ $:=f$ $\text{ for }f\in[0,1]$ (3)
$u(\phi_{1}\land^{p}\phi_{2})$ $:=\overline{\mu}_{p}(u(\phi_{1}),u(\phi_{2}))$ $\text{for }\phi_{1},\phi_{2}:\text{{FPL}}$ (4)
$u(\neg\phi)$ $:=1-u(\phi)$ $\text{for }\phi:\text{{FPL}}$ (5)
$u(\phi_{1}\lor^{p}\phi_{2})$ $:=u(\neg(\neg\phi_{1}\land^{p}\neg\phi_{2}))$ $\text{for }\phi_{1},\phi_{2}:\text{{FPL}}$ (6)
$u([\phi]_{\delta})$ $:=\frac{u(\phi)+\max(\delta,0)}{1+\delta}$ $\text{for }\phi:\text{{FPL}}$ (7)
These semantics preserve logical relationships between objectives. The conjunction $\land^{p}$ combines objectives that must be fulfilled together, with $p\leq 0$ controlling strictness (formalized in Definition IV-G). The disjunction $\lor^{p}$, defined via De Morgan’s laws, allows focusing on the objective with highest marginal benefit. The offset $\phi]_{\delta}$ creates priority by raising the baseline fulfillment of $\phi$, implementing a form of lexicographic ordering where objectives with lower offsets are prioritized. By making fulfillment composition explicit, FPL allows practitioners to seperate concerns and reason about the trade-offs between objectives in a principled way. We expect practitioners to first define fulfillment values for their objectives, testing whether their intention for the fulfillment of objectives maps well on to [0,1], and then use FPL to compose them in a way that is consistent with their priorities.
@@ IV-C Relation to Fuzzy Logic
FPL similar to fuzzy logic, generalizes boolean operations to the continuous domain $[0,1]$, meaning, at the limits $\{0,1\}$ it is equivalent to boolean logic. Choosing $p=0$, our conjunction operator becomes the geometric mean ($\forall_{x,y\in[0,1]},x\land^{0}y=\sqrt{x\cdot y}$) closely resembling the product t-norm [41] ($\forall_{x,y\in[0,1]},x\land y=x\cdot y$) commonly used in fuzzy logic. Also, when $\to-\infty, p\land^{p}$ becomes the minimum operator, which is equivalent to the fuzzy logic’s minimum t-norm.
@@ IV-E Toy Example
Priority Offset (${\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}f_{0}}]_{\delta}\land^{p}{\color[rgb]{0.1171875,0.56640625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.56640625,1}f_{1}}$) [Uncaptioned image]: The offset gives precedence to $\color[rgb]{0.1171875,0.56640625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.56640625,1}f_{1}$} until it reaches the threshold established by $\delta$, at which point $\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}f_{0}$} becomes important as well. This priority-based curriculum allows for coherent optimization, as opposed to methods that change the reward function requiring Q-values to be re-trained.
@@ IV-G Guarantees in FPL
This bound guarantees that when a power mean outputs value $y$, every input component must have at least fulfillment $\sqrt[{}^{p}]{n(y^{p}-1)+1}$.
@@
This bound is crucial for practitioners providing concrete guarantees about objective fulfillment. For example, with two objectives ($n=2$) and p=2p=-2, achieving an output of 0.9 guarantees each individual objective has at least 0.38 fulfillment. This minimum-fulfillment is particularly important in safety-critical applications where we need to ensure no objective is severely underperforming. The parameter $p$ allows practitioners to trade off between stronger guarantees (more negative $p$) and easier optimization (less negative $p$), as stricter bounds require more precise balancing of objectives.
@@
where $\mathbbm{1}_{i}$ denotes a vector of zeros with a 1 in position ii. This states that for any change $\delta$ to component $i$, there exists a change $\delta^{\prime}$ to component $j$ that maintains the same power mean.
@@
For any vector $\vec{\mathbf{f}}$, there exists a vector with n1n-1 ones and a value $v$ that has the same power mean but with $v$ being bounded from above by the minimum of $\vec{\mathbf{f}}$.
@@
For a vector with n1n-1 ones and one value $f$, if $\overline{\mu}_{p}((\vec{\mathbf{1}}_{n-1},f))=y$ then:
@@ V-A Mitigating Overestimation Bias
[…]
@@
   $L^{\text{TD}}\leftarrow\overline{\mu}_{2}(\vec{\mathbf{y}}^{\text{TD}}$- $\mathbf{FQ}$$(s,a))$
@@ VI-C1 Pendulum-v1
$\theta^{2}-0.1\,\dot{\theta}^{2}-0.001\,\text{torque}^{2}$ $F_{\text{angle}}^{2}\land^{p}F_{\text{actuation}}$
@@ VI-C2 Reacher-v4
[…]
@@ VI-C3 Hopper-v4
[…]
@@ VI-C4 LunarLanderContinuous-v2
$\text{distance\_reward}+\text{velocity\_reward}+\text{angle\_reward}$
$+10\cdot\text{legs\_contact}-0.3\cdot|main\_engine|$
$-0.03\cdot|side\_engines|+\text{terminal\_reward}$
@@
Here the offsets create a natural curriculum during training: the agent first focuses on basic proximity ($F_{\text{near}}$), then simultaneously addresses precise positioning, leg contact, and landing ($F_{\text{very\_near}}]_{0.1}$, $F_{\text{legs}}]_{0.1}$, $F_{\text{landed}}]_{0.1}$), and finally optimizes fuel efficiency ($F_{\text{fuel}}]_{0.5}$) once the primary landing objectives are reasonably satisfied. The conjunction ensures all objectives must ultimately be satisfied for successful landing.
@@ VI-D Behavioral Analysis
A note on parameter selection in FPL: FPL is robust to reasonable variations in power mean parameters and offsets, we choose $p$ as either 0 or 1-1, which primarily serve to optimize sample efficiency rather than fundamentally changing the desired behavior. For example, not squaring the angle term in Pendulum would still result in an upright pendulum, but with slower convergence due to more conservative actions. This behavioral consistency persists across training runs, unlike linear weighted reward functions that often converge to different local optima depending on initialization.